OpenStudy (anonymous):

a sports car accelerates from rest to 95km/h in 6.2 s. What is the average acceleration in m/s^2

OpenStudy (aaronq):

\(\Delta a=\dfrac{\Delta v}{\Delta s}=\dfrac{v_f-v_i}{t_2-t_1}=\dfrac{95~km/h-0~km/h}{6.2s-0s}=\dfrac{95~km/h}{6.5~s}\)

OpenStudy (aaronq):

i forgot to convert km/h to m/s, hold on

OpenStudy (aaronq):

\(95~km/h*\dfrac{1000~m}{1km}*\dfrac{1h}{60 min}*\dfrac{1min}{60s}=26.4 m/s\) so the above should be \(\Delta a=\dfrac{26.4 m/s}{6.5~s}\)